OpenStudy (anonymous):

Solve for m: w = m - 5L. Solve for m: w = m - 5L. @Mathematics

OpenStudy (anonymous):

m=w+5L, isnt it?

OpenStudy (anonymous):

\[\Large \begin{array}{l} w = m - 5l\\ w + 5l = m - 5l + 5l\\ w + 5l = m \end{array}\]